OpenStudy (anonymous):

If 5p = -100 + 3Q^2, which of the following could Q possibly be?

OpenStudy (anonymous):

Q = square root of 5 times p minus 100 over 3 Q = square root of 5 times p plus 100 over 3 Q = square root of negative 5 times p plus 100 over 3 Q = square root of negative 5 times p minus 100 over 3

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

1.add 100 to both sides and then divide by 3 2. take the squareroot

OpenStudy (anonymous):

so that would be q^2 = 100+5p/3

OpenStudy (anonymous):

the sqrt of 100 + 5p/3 = q

OpenStudy (anonymous):

So it would be the second one?

OpenStudy (whpalmer4):

if you write fractions like that, you need to use parentheses! \[5p = -100+3Q^2\]\[5p+100 = -100+100 + 3Q^2\]\[(5p+100)/3 = Q^2\] \[Q = \sqrt{(5p+100)/3}\]

OpenStudy (anonymous):

\[3 Q^2=5P+100\] \[Q^2=\frac{ 5P+100 }{ 3 }\] Q=?

OpenStudy (whpalmer4):

100 + 5p/3 is not the same as \[\frac{100+5p}{3}\]

OpenStudy (anonymous):

oh ok thanks you both so much for the help

OpenStudy (whpalmer4):

yw